# Edge Delayed Deep Deterministic Policy Gradient: efficient continuous control for edge scenarios

### v1

@@ A.2 Hyper-parameters
In table 4, we report the hyperparameters used for the simulations. For a fair comparison, the hyper-parameters that could lead to an unfair setting, such as the stepsize, have been kept constant throughout all the methods. For the hyper-parameters that were not common to all of them, we used the one reported in the respective original papers. Regarding the proposed methods, the only parameters that have been varied are the hyperparameters for controlling the trade-off between overestimation and underestimation, formerly $\alpha,\beta$. However, to avoid cherrypicking of such parameters, only good guesses have been used, and are all reported in fig. 6. Regarding the decay, we observed little improvement in using both a linear decay and an exponential decay during the execution. Since such decay would be part of non-trivially tunable hyper-parameters, and we wanted to keep the method as simple as possible, we decided to use $\lambda(t)=1$, so no decay has been applied during any of the training reported throughout the paper. Thus, all the curves report learning done with fixed $\alpha,\beta$. Regarding the noise distribution used for the action in the target estimation, we used the clipped Normal distribution introduced in [5].
Table 4: List of hyperparameters used for training.
Hyper-parameter TD3 DDPG EdgeDDPG EdgeD3 SAC
Critic learning-rate $0.0003$ $0.0003$ $0.0003$ $0.0003$ $0.0003$
Actor learning-rate $0.0003$ $0.0003$ $0.0003$ $0.0003$ $0.0003$
Optimizer Adam Adam Adam Adam Adam
Target Update Rate ($\tau_{1},\tau_{2}$) $0.005$ $0.005$ $0.005$ $0.005$ $0.005$
Batch-size $256$ $256$ $256$ $256$ $256$
Training iteration per step $1$ $1$ $1$ $1$ $1$
Discount factor $0.99$ $0.99$ $0.99$ $0.99$ $0.99$
Exploration policy $N(0,0.2)$ $N(0,0.2)$ $N(0,0.2)$ $N(0,0.2)$ learnt
Entropy - - - - 0.5
Actor heads count $1$ $1$ $1$ $1$ $1$
Actor update delay ($d$) $2$ $1$ $1$ $2$ $1$
@@ Appendix B Ablation over various tradeoffs
Figure 6: Comparing EdgeD3 with EdgeDDPG in continuous control tasks. Plots are from 10 random seeds for simulator and network initializations, smoothed for visualization. Evaluations of Return are performed every 5000 time steps, plots show mean and half a standard deviation, over 10 episodes.
@@ Appendix C Comparison with environment steps
Figure 7: Comparing EdgeD3 with baselines in continuous control tasks using environment timesteps. Plots are from 10 random seeds for simulator and network initializations, smoothed for visualization. Evaluations of Return are performed every 5000 time steps, plots show mean and half a standard deviation, over 10 episodes.

### v2

@@ A-B Hyper-parameters and further ablations
In table IV, we report the hyperparameters used for the simulations. For a fair comparison, the hyper-parameters that could lead to an unfair setting, such as the stepsize, have been kept constant throughout all the methods. For the hyper-parameters that were not common to all of them, we used the one reported in the respective original papers. Regarding the proposed methods, the only parameters that have been varied are the hyperparameters for controlling the trade-off between overestimation and underestimation, formerly $\alpha,\beta$. However, to avoid cherrypicking of such parameters, only good guesses have been used, and are all reported in fig. 8. Regarding the decay, we observed little improvement in using both a linear decay and an exponential decay during the execution. Since such decay would be part of non-trivially tunable hyperparameters, and we wanted to keep the method as simple as possible, we decided to use $\lambda(t)=1$, so no decay has been applied during any of the training reported throughout the paper. Thus, all the curves report learning done with fixed $\alpha,\beta$. Regarding the noise distribution used for the action in the target estimation, we used the clipped Normal distribution introduced in [14].
TABLE IV: List of hyperparameters used for training.
Hyper-parameter TD3 DDPG EdgeDDPG EdgeD3 SAC PPO
Critic learning-rate $0.0003$ $0.0003$ $0.0003$ $0.0003$ $0.0003$ $0.0003$
Actor learning-rate $0.0003$ $0.0003$ $0.0003$ $0.0003$ $0.0003$ $0.0003$
Optimizer Adam Adam Adam Adam Adam Adam
Target Update Rate ($\tau_{1},\tau_{2}$) $0.005$ $0.005$ $0.005$ $0.005$ $0.005$ -
Batch-size $256$ $256$ $256$ $256$ $256$ $256$
Training iteration per step 11 11 11 11 11 $10$
Discount factor $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$
Exploration policy $N(0,0.2)$ $N(0,0.2)$ $N(0,0.2)$ $N(0,0.2)$ learnt -
Entropy - - - - 0.5 -
Actor heads count 11 11 11 11 11 -
Actor update delay ($d$) 22 11 11 22 11 -
A-C Comparison to Actor Critic
Another branch of algorithms in reinforcement learning is the family of actor-critic methods. Actor-critic algorithms combine the advantages of value-based and policy-based approaches by maintaining both a parameterized policy function (the actor) and a value function (the critic). This structure enables the actor to update its policy in the direction suggested by the critic, which estimates the expected return. Classical actor-critic methods, however, are often susceptible to issues such as high variance in policy gradient estimates and instability during training. Yet, they offer learning policies with lower memory requirements than DDPG-like algorithms, as they require only 2 networks, the critic and the actor, without target networks. In this section, we thus compare our proposed approach to Proximal Policy Optimization [56], which represents the state-of-the-art algorithm in this family, and the most adopted one, thanks to its performance and robustness.
Regarding memory consumption, we find that PPO utilizes 18% less RAM thanks to the fact that no target network is required. Yet, given that the target network is not optimized by gradient descent, the memory consumption required to store it is less than expected (1 out of 3 neural networks). However, PPO, in order to obtain its performance, utilized multiple steps of updates, with an off-policy loss. These multiple updates makes PPO much more expensive than the proposed approach, requiring 35% more compute for the same training budget. The overall performances are shown in fig. 7. Regarding EdgeD3, the tradeoff $\alpha=1,\beta=2$ has been used for the comparison, as per the final comparison in fig. 4. Regarding PPO, as reported in the original paper, 1 step of updates was done for the value network, and 10 for the actor. Given that such an update count increases the effectiveness of PPO while also increasing computational cost, we did not find a major improvement in increasing over 10 such update count.
Figure 7: Comparing EdgeD3 with PPO and TD3 in continuous control tasks. Plots are from 10 random seeds for simulator and network initializations, smoothed for visualization. Evaluations of Return are performed every 5000 time steps, plots show mean and standard deviation, over 10 episodes.
@@ Appendix B Ablation over various tradeoffs
Figure 8: Comparing EdgeD3 with EdgeDDPG in continuous control tasks. Plots are from 10 random seeds for simulator and network initializations, smoothed for visualization. Evaluations of Return are performed every 5000 time steps, plots show mean and standard deviation, over 10 episodes.
Figure 9: Comparing EdgeD3 with baselines in continuous control tasks using environment timesteps. Plots are from 10 random seeds for simulator and network initializations, smoothed for visualization. Evaluations of Return are performed every 5000 time steps, plots show mean and standard deviation, over 10 episodes.
@@ Appendix C Comparison with environment steps
C-A Networks architectures
For TD3, DDPG, EdgeDDPG, PPO, and EdgeD3, the actor consists of two fully connected layers of 256 units each with ReLU activations, followed by a fully connected output layer mapping to the action dimension with a hyperbolic tangent activation. For SAC, the actor has the same structure except that the output layer maps to twice the action dimension. For all methods, the critic consists of two fully connected layers of 256 units each with ReLU activations, followed by a single-unit output layer.